How do u find in two sets of equations have points in common
2x^2+3y^2=19
x^2+y^2=9

Answer by Alan3354(69443)

How do u find in two sets of equations have points in common
2x^2+3y^2=19
x^2+y^2=9 --> y^2 = 9-x^2
sub for y^2 in the 1st eqn and solve for x
2x^2+3y^2=19
2x^2 + 3(9-x^2) = 19
-x^2 = -8
x = +/- sqrt(8)
The solutions are real numbers, so there are intersections.
Sub each value for x and find 2 values of y (in either equation).
4 intersections.